### v1

# Decoupling and Damping: Structurally-Regularized Gradient Matching for Multimodal Graph Condensation

Abstract.
In critical web applications such as e-commerce and recommendation systems, multimodal graphs integrating rich visual and textual attributes are increasingly central, yet their large scale introduces substantial computational burdens for training Graph Neural Networks (GNNs). While Graph Condensation (GC) offers a promising solution by synthesizing smaller datasets, existing methods falter in the multimodal setting. We identify a dual challenge causing this failure: (1) conflicting gradients arising from semantic misalignments between modalities, and (2) the GNN’s message-passing architecture pathologically amplifying this gradient noise across the graph structure. To address this, we propose Structurally-Regularized Gradient Matching (SR-GM), a novel condensation framework tailored for multimodal graphs. SR-GM introduces two synergistic components: first, a gradient decoupling mechanism that resolves inter-modality conflicts at their source via orthogonal projection; and second, a structural damping regularizer that acts directly on the gradient field. By leveraging the graph’s Dirichlet energy, this regularizer transforms the topology from a noise amplifier into a stabilizing force during optimization. Extensive experiments demonstrate that SR-GM significantly improves accuracy and accelerates convergence compared to baseline methods. Ablation studies confirm that addressing both gradient conflict and structural amplification in tandem is essential for achieving superior performance. Moreover, the condensed multimodal graphs exhibit strong cross-architecture generalization and promise to accelerate applications like Neural Architecture Search. This research provides a scalable methodology for multimodal graph-based learning in resource-constrained environments.
Multimodal Graph Condensation, Decoupled Gradient Matching, Structural Regularization
$c$s: Information systems Data miningccs: Theory of computation Graph algorithms analysis
1. Introduction
The deep integration of the Internet and multimedia technologies has spawned a massive amount of graph-structured data containing rich heterogeneous information. From users in social networks and the multimedia content they publish (khattar2019mvae; wang2012multimodal; su2021detecting; tao2020mgat; safavi2020codex) to product graphs associated with text descriptions and product images on e-commerce platforms (xu2020product; deng2023construction), these multimodal graphs have become the core of modern web applications. Graph neural networks (GNNs) (kipf2016semi; rong2019dropedge; zhang2018link), as powerful tools for processing such data, have achieved remarkable results in tasks such as node classification and link prediction. However, their success relies heavily on large-scale datasets (safavi2020codex; wan2023scalable), which leads to high computational costs, including lengthy training time, huge memory consumption, and complex hyperparameter tuning (zhang2022efficient; cai2024multimodal).
To address this challenge, a data-centric approach, dataset condensation (wang2018dataset; zhao2020dataset), has emerged. Its goal is to refine the original large-scale dataset into a small synthetic dataset with highly concentrated information, so that the model trained on this synthetic dataset can achieve performance comparable to that trained on the full dataset. In the field of computer vision, gradient matching (zhao2020dataset) has become a benchmark method for achieving this goal. It efficiently learns highly representative synthetic samples by matching the model’s training gradient trajectories on real data and synthetic data. Recently, researchers have begun to extend this paradigm to graph-structured data and proposed a variety of graph condensation methods (jin2021graph; gao2023multiple; liu2022graph). These methods are usually formulated as a complex bi-level optimization problem, aiming to learn a miniature synthetic graph structure and its node features through objectives such as gradient matching (jin2021graph) or distribution matching (liu2022graph).
Although these methods show potential on unimodal graphs (with node features from a unimodal source), their performance degrades significantly when directly applied to more general multimodal graphs (where nodes have multiple modal features such as text and images). Through in-depth analysis, we find that the root cause lies in the conflict between multimodal gradient signals and the pathological amplification effect of the graph structure on this conflict. Specifically, there may be a semantic gap between different modalities (such as text and image) (liang2022mind; zhang2025beyond), which leads to the generation of gradients in opposite directions during the optimization process. More importantly, the inherent message passing mechanism of GNN will quickly propagate and amplify the gradient conflicts of these local nodes to the entire graph, thus making the optimization process of the condensation algorithm based on gradient matching unstable and ineffective.
To overcome this problem, this paper proposes a novel structurally regularized gradient matching (SR-GM) framework designed for multimodal graph condensation. The core ideas of SR-GM are twofold: (1) Gradient decoupling: At its core, we adopt the idea of conflict disentanglement from multi-task learning (yu2020gradient; chai2024towards). Specifically, for the conflicting gradients arising from different modalities on a synthetic node, we apply orthogonal projection to eliminate their inherent contradictions, thereby achieving synergistic inter-modal gradient collaboration; (2) Structural damping: On the propagation path, we innovatively introduce a graph Laplacian regularization term that acts on the gradient field rather than node features. This regularization term forces the optimization directions of adjacent nodes to be consistent during the condensation process, thereby transforming the graph structure from a ”conflict amplifier” to an ”optimization damper”, effectively suppressing the spread of gradient noise.
We conducted extensive experiments on multiple multimodal graph benchmark datasets. Our results demonstrate that SR-GM significantly outperforms existing graph condensation baselines. Detailed ablation studies confirm the necessity and effectiveness of the two components, gradient decoupling and structural damping. Furthermore, the condensed graphs generated by SR-GM exhibit excellent generalization across diverse GNN architectures, highlighting its robustness and practicality.
The main contributions of this paper are as follows:
  • We provide the first in-depth analysis of the mechanisms of gradient conflict in multimodal graph condensation and its negative effects amplified by graph structure, providing a novel theoretical perspective for understanding the underlying problem.
  • We propose the SR-GM method, which systematically addresses the core challenges of multimodal graph condensation through the dual design of gradient decoupling and structural damping.
  • We validate the superior performance of SR-GM through rigorous experiments and demonstrate the strong generalization of the generated condensed data in cross-architecture evaluations.
2. Related Work
2.1. Graph Dataset Condensation
Dataset Condensation (DC), also known as Dataset Distillation, aims to synthesize a small, informative dataset from a large original one such that models trained on the synthetic set achieve performance comparable to those trained on the original data. This paradigm was first pioneered in the image domain, where methods were developed to optimize synthetic pixels directly through a bi-level optimization framework (wang2018dataset). The core idea is to match certain properties between the real and synthetic data. Various matching criteria have been proposed, including matching model performance (zhao2020dataset) and feature distributions (zhao2023dataset; wang2022cafe). Among these, matching the gradients of model parameters has been proven to be a particularly effective strategy (zhao2020dataset). The extension of these techniques to the graph domain, termed Graph Condensation (GC), is a nascent and challenging endeavor. Jin et al. (jin2021graph) first extended DC to graphs, introducing a node-level condensation method. Unlike image data, graph data involves an intricate coupling between node attributes and topological structure, making the joint synthesis of features and edges a complex optimization task. Current GC methods primarily focus on preserving information in single-modal graphs, exploring strategies such as gradient matching (jin2021graph), training trajectory matching (zheng2023structure), and distribution matching (liu2022graph). Recent research has evolved from developing efficient and interpretable graph condensation methods (fang2024exgc; wang2024fast) to creating specialized techniques tailored for specific graph types, such as heterogeneous graphs and signed graphs (gao2024heterogeneous; li2025structure). However, these methods largely overlook the unique challenges posed by multimodal graphs, where node features themselves are composed of heterogeneous information from sources like text and images. The problem of inter-modal signal conflict within the GC framework remains a critical and underexplored gap. Our work aims to bridge this gap.
2.2. Multimodal Graph Learning
Multimodal Graph Learning (MGL) has emerged as a vital field for analyzing complex real-world systems where entities are described by multiple data types (e.g., text, images, audio). MGL aims to leverage the relational structure of graphs to fully explore associations both within and between different modalities. This is particularly challenging because data fusion must be guided by the graph topology, which itself may vary across modalities. A key challenge in MGL is bridging the ”semantic gap”, where different modalities may provide conflicting or inconsistent representations of the same entity (zhu2025mosaic; liang2022mind). For instance, two products might be visually similar but have different textual descriptions, creating discrepancies that unimodal models cannot resolve. To address this, significant research has focused on feature alignment, which aims to map representations from different modalities into a unified embedding space. Contrastive learning frameworks, such as CLIP (radford2021learning), have been notably successful in learning aligned vision-language representations. From an architectural perspective, methods like Multimodal Graph Convolutional Networks (MGCN) (wei2019mmgcn) and Multimodal Graph Attention Networks (MGAT) (tao2020mgat) have been proposed to explicitly model and fuse information from different modalities. While these MGL methods focus on designing more sophisticated Graph Neural Network (GNN) architectures to better handle existing multimodal data, they do not address the data-centric challenge of scalability (ninggraph4mm; cai2024multimodal). The question of how to effectively compress large-scale multimodal graphs, particularly in a way that handles the feature conflicts inherent in the condensation process, remains largely unaddressed. Our work bridges this gap by introducing a condensation method specifically designed with an optimization scheme attuned to the uniqueness of multimodal graphs.
[…]
3. Methodology
3.1. Problem Formulation
We consider the problem of multimodal graph condensation, where the objective is to distill a large-scale multimodal graph into a compact yet informative synthetic graph while preserving the essential information for downstream tasks.
Let the large-scale original graph be denoted as $\mathcal{T}=(\mathbf{X},\mathbf{A},\mathbf{Y})$, where $\mathbf{X}\in\mathbb{R}^{N\times d}$ is the node feature matrix for $N$ nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix, and $\mathbf{Y}\in\{0,1,\dots,C-1\}^{N}$ are the node labels over $C$ classes. In the context of multimodal graphs, the $d$-dimensional feature vector for each node is a concatenation of features from different modalities, i.e., $\mathbf{X}_{i}=[\mathbf{x}_{i,\text{text}}\oplus\mathbf{x}_{i,\text{image}}]$.
The objective of graph condensation is to synthesize a small, information-rich graph $\mathcal{S}=(\mathbf{X}_{\mathcal{S}},\mathbf{A}_{\mathcal{S}},\mathbf{Y}_{\mathcal{S}})$ with $N^{\prime}$ nodes ($N^{\prime}\ll N$), such that a Graph Neural Network (GNN) model, $\phi_{\theta}$, trained on $\mathcal{S}$ achieves comparable performance to the same model trained on $\mathcal{T}$.
3.2. Gradient Matching for Multimodal Graph
This goal is formally captured as a bi-level optimization problem (jin2022condensing). The outer loop seeks to find the optimal synthetic graph $\mathcal{S}^{*}$ by minimizing a loss $\mathcal{L}_{\text{outer}}$ that evaluates a model trained on $\mathcal{S}$. The inner loop trains this model on $\mathcal{S}$ to convergence to obtain the optimal parameters $\theta_{\mathcal{S}}^{*}$. This can be written as:
(1) $\mathcal{S}^{*}=\arg\min_{\mathcal{S}}\mathcal{L}_{\text{outer}}(\phi_{\theta_{\mathcal{S}}^{*}},\mathcal{T})\quad\text{s.t.}\quad\theta_{\mathcal{S}}^{*}=\arg\min_{\theta}\mathcal{L}_{\text{inner}}(\phi_{\theta},\mathcal{S})$
Solving this formulation directly is computationally prohibitive, as it requires fully training a GNN in the inner loop for every single update to the synthetic graph $\mathcal{S}$ in the outer loop.
The Gradient Matching (GM) paradigm offers an efficient and effective approximation to this problem (jin2021graph). Instead of matching the performance of fully converged models, GM aims to align the learning dynamics by matching the parameter gradients at each step of a single training trajectory. The process involves two interleaved optimizations: an outer-loop optimization for the synthetic data $\mathcal{S}$ and an inner-loop update for the model parameters $\theta$. The objective is to minimize the cumulative distance between the gradients over the entire trajectory:
(2) $\min_{\mathcal{S}}\mathbb{E}_{\theta_{0}\sim P_{\theta}}\left[\sum_{t=0}^{T-1}D(\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t}),\nabla_{\theta}\mathcal{L}_{\mathcal{T}}(\theta_{t}))\right]$
where the model parameters are updated in an inner loop according to the synthetic data: $\theta_{t+1}=\theta_{t}-\eta_{\theta}\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t})$. Here, $\mathcal{L}_{\mathcal{S}}$ and $\mathcal{L}_{\mathcal{T}}$ are the task losses, and $D(\cdot,\cdot)$ is a distance metric between the gradient vectors.
Traditional gradient matching methods, while successful in unimodal domains, face significant challenges when applied to multimodal graphs due to two fundamental issues.
3.3. Inter-Modality Gradient Conflict
The outer loop of the GM algorithm is tasked with updating the synthetic graph $\mathcal{S}$ to minimize the distance to a target gradient, $\mathbf{g}_{\mathcal{T}}=\nabla_{\theta}\mathcal{L}_{\mathcal{T}}$, sampled from the real data. In the multimodal setting, this target gradient becomes inherently unreliable.
To formalize this, we frame the GNN’s learning process as a Multi-Task Learning (MTL) problem (liu2021conflict; wang2020gradient), where each modality constitutes a distinct task being learned by the shared backbone, parameterized by $\theta_{\text{sh}}$. The total gradient $\mathbf{g}_{\mathcal{T}}$ is a linear combination of the gradients from each modality:
(3) $\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{total}}=\sum_{k\in\{\text{text, image}\}}w_{k}\nabla_{\theta_{\text{sh}}}\mathcal{L}_{k}$
This simple summation becomes problematic due to the feature-level ”semantic gap” inherent in multimodal data (feng2023mkvse; liu2024rng). This misalignment generates opposing update signals, leading to a gradient conflict(mahapatra2020multi; ma2020efficient), formally defined as the condition where the gradients of two tasks have a negative cosine similarity:
(4) $\langle\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{text}},\nabla_{\theta_{\text{sh}}}\mathcal{L}_{\text{image}}\rangle<0$
When this occurs, the target gradient $\mathbf{g}_{\mathcal{T}}$ becomes a high-variance and suboptimal signal. It represents a blind compromise between conflicting objectives rather than a stable, Pareto-optimal update direction (mahapatra2020multi). Consequently, the outer loop of GM is tasked with matching a noisy, erratically moving target, which fundamentally destabilizes the meta-optimization of $\mathcal{S}$.
3.4. Structural Amplification of Gradient Noise
The instability is further compounded by the GNN’s message-passing mechanism (feng2022powerful), which corrupts the update step of the outer loop itself. The outer loop updates the synthetic data $\mathcal{S}$ using a meta-gradient, computed by differentiating the matching loss with respect to $\mathcal{S}$. This requires backpropagating through the computation of the synthetic gradient $\mathbf{g}_{\mathcal{S}}$, and therefore, through the GNN architecture.
During this meta-gradient computation, the gradient of the loss with respect to a node’s features is influenced by the gradients of its neighbors, causing the gradient to flow backward along the graph edges:
(5) $\nabla_{h_{i}^{(k-1)}}\mathcal{L}=\sum_{j\mid i\in\mathcal{N}(j)}\frac{\partial\mathcal{L}}{\partial h_{j}^{(k)}}\frac{\partial h_{j}^{(k)}}{\partial h_{i}^{(k-1)}}$
This dynamic transforms the GNN’s computational graph into a medium for noise propagation. The localized gradient conflicts at individual nodes are not contained; instead, they are amplified and propagated throughout the graph structure. This creates a highly non-smooth gradient field over the synthetic graph, where the optimization directions for adjacent nodes can be wildly different.
This structural amplification can be formally quantified using the Dirichlet Energy (zhou2021dirichlet) of the feature gradient field. Let $\mathbf{G}_{\mathcal{S}}$ be the $N^{\prime}\times d$ matrix where each row is the feature gradient $\nabla_{\mathbf{X}_{\mathcal{S},i}}\mathcal{L}$ for a synthetic node $i$, and let $\mathbf{L}_{\mathcal{S}}$ be the graph Laplacian. The Dirichlet energy of the gradient field is:
(6) $E(\mathbf{G}_{\mathcal{S}})$ $=\mathrm{tr}(\mathbf{G}_{\mathcal{S}}^{T}\mathbf{L}_{\mathcal{S}}\mathbf{G}_{\mathcal{S}})$
(7) $=\frac{1}{2}\sum_{(i,j)\in E_{\mathcal{S}}}w_{ij}\|\nabla_{\mathbf{X}_{\mathcal{S},i}}\mathcal{L}-\nabla_{\mathbf{X}_{\mathcal{S},j}}\mathcal{L}\|_{2}^{2}$
A high Dirichlet energy signifies a chaotic gradient field. The structural amplification mechanism ensures this energy remains high, creating a rugged optimization landscape for $\mathcal{S}$ and causing the meta-gradient to be unstable. This persistent instability in the gradient field directly undermines the condensation objective, making it difficult to learn a small yet representative synthetic graph.
3.5. The Proposed SR-GM Framework
To address this dual challenge, as shown in Figure 1, we propose Structurally-Regularized Gradient Matching (SR-GM), a framework that re-engineers the condensation process through two synergistic components: Gradient Decoupling and Structural Damping.
3.5.1. Parametrization of the Synthetic Graph
In SR-GM, the synthetic graph $\mathcal{S}$ is parameterized by its node features $\mathbf{X}_{\mathcal{S}}$ and the parameters $\Phi$ of a structure generator network $g_{\Phi}$. The node features $\mathbf{X}_{\mathcal{S}}$ are learned directly. Following GCond (jin2021graph), the adjacency matrix $\mathbf{A}_{\mathcal{S}}$ is not directly parameterized but is generated as a function of the node features: $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$.
Specifically, the weight of an edge between nodes $i$ and $j$ is computed by a shared MLP, which is a multi-layer perceptron with parameters $\Phi$:
(8) $(\mathbf{A}_{\mathcal{S}})_{ij}=\text{Sigmoid}\left(\frac{\text{MLP}_{\Phi}([\mathbf{x}_{i};\mathbf{x}_{j}])+\text{MLP}_{\Phi}([\mathbf{x}_{j};\mathbf{x}_{i}])}{2}\right)$
where $\cdot$$\cdot$] denotes concatenation. This formulation makes the graph structure differentiable with respect to both the node features $\mathbf{X}_{\mathcal{S}}$ and the generator parameters $\Phi$, and naturally enforces symmetry for undirected graphs. The set of learnable parameters for the outer loop is thus $\{\mathbf{X}_{\mathcal{S}},\Phi\}$.
3.5.2. Gradient Decoupling via Orthogonal Projection
To resolve the root cause of instability, we perform ”gradient surgery” on the feature gradients (liu2021conflict; yu2020gradient) of each synthetic node. This process is a prerequisite for computing the structural regularization term and is integrated into the outer loop’s meta-gradient calculation.
Working Mechanism: At each inner-loop step $t$, after computing the synthetic loss $\mathcal{L}_{\mathcal{S}}(\theta_{t})$, we first compute the gradient of this loss with respect to the synthetic features $\mathbf{X}_{\mathcal{S}}$. This yields the feature gradient matrix $\mathbf{G}_{\mathcal{S}}\in\mathbb{R}^{N^{\prime}\times d}$:
(9) $\mathbf{G}_{\mathcal{S}}=\nabla_{\mathbf{X}_{\mathcal{S}}}\mathcal{L}_{\mathcal{S}}(\theta_{t})$
Since the synthetic features $\mathbf{X}_{\mathcal{S}}$ are a concatenation of modal features, the resulting gradient matrix $\mathbf{G}_{\mathcal{S}}$ has the same structure. For each node $s_{i}$, we can partition its corresponding row vector in $\mathbf{G}_{\mathcal{S}}$ to obtain the modality-specific gradients, $\mathbf{g}_{i,\text{text}}$ and $\mathbf{g}_{i,\text{image}}$.
If a conflict exists ($\langle\mathbf{g}_{i,\text{text}},\mathbf{g}_{i,\text{image}}\rangle<0$), we project each gradient onto the normal plane of the other to eliminate the conflicting components, following the PCGrad (yu2020gradient) methodology:
(10) $\mathbf{g^{\prime}}_{i,\text{text}}$ $=\mathbf{g}_{i,\text{text}}-\frac{\langle\mathbf{g}_{i,\text{text}},\mathbf{g}_{i,\text{image}}\rangle}{\|\mathbf{g}_{i,\text{image}}\|^{2}}\mathbf{g}_{i,\text{image}}$
(11) $\mathbf{g^{\prime}}_{i,\text{image}}$ $=\mathbf{g}_{i,\text{image}}-\frac{\langle\mathbf{g}_{i,\text{image}},\mathbf{g}_{i,\text{text}}\rangle}{\|\mathbf{g}_{i,\text{text}}\|^{2}}\mathbf{g}_{i,\text{text}}$
This operation yields a decoupled feature gradient matrix $\mathbf{G^{\prime}}_{\mathcal{S}}$, where each row is the reconciled gradient $\mathbf{g^{\prime}}_{i}=\mathbf{g^{\prime}}_{i,\text{text}}+\mathbf{g^{\prime}}_{i,\text{image}}$.
3.5.3. Structural Damping via Gradient Field Regularization
To counteract the structural amplification of noise, we introduce a novel regularization term that operates directly on the decoupled gradient field. Unlike traditional graph regularization (yang2021rethinking) on node features, which GNNs already implicitly perform, our regularizer enforces smoothness on the optimization directions themselves. Let $\mathbf{G^{\prime}}_{\mathcal{S}}$ be the matrix of decoupled feature gradients from Equation (7). The structural regularization term is the Dirichlet energy of this gradient field:
(12) $\mathcal{R}_{\text{struct}}(\mathbf{G^{\prime}}_{\mathcal{S}})=\mathrm{tr}(\mathbf{G^{\prime}}_{\mathcal{S}}^{\top}\mathbf{L}_{\mathcal{S}}\mathbf{G^{\prime}}_{\mathcal{S}})=\sum_{(i,j)\in E_{\mathcal{S}}}w_{ij}\|\mathbf{g^{\prime}}_{i}-\mathbf{g^{\prime}}_{j}\|^{2}$
This term penalizes large differences in optimization directions between connected nodes, forcing the gradient field to be smooth. It thereby transforms the graph structure from a ”conflict amplifier” into an ”optimization damper.”
3.5.4. Final Objective and Algorithm
By integrating both components, the final SR-GM objective function becomes:
(13) $\min_{\mathcal{S}}\mathbb{E}_{\theta_{t}}\left[D(\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t}),\nabla_{\theta}\mathcal{L}_{\mathcal{T}}(\theta_{t}))+\lambda\cdot\mathcal{R}_{\text{struct}}\right]$
where $\lambda$ is a hyperparameter balancing gradient matching accuracy with gradient field smoothness.The gradient matching loss $\mathcal{L}_{\text{gm}}=D\left(\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t}),\nabla_{\theta}\mathcal{L}_{\mathcal{T}}(\theta_{t})\right)$ follows the implementation in GCond (jin2021graph). SR-GM design fundamentally repurposes the graph structure from a source of instability into a tool for stabilization. The complete procedure, which correctly interleaves the outer-loop update of $\mathcal{S}$ and the inner-loop update of $\theta$, is detailed in Algorithm 1.
Algorithm 1 Structure-Regularized Gradient Matching (SR-GM)
1:Original graph $\mathcal{T}$, GNN model $\phi_{\theta}$, hyperparameters $\lambda$, $\eta_{\mathcal{S}}$, $\eta_{\Phi}$, $\eta_{\theta}$, outer-loop iterations $K$, inner-loop iterations $T$
2:Optimized synthetic features $\mathbf{X}_{\mathcal{S}}$ and structure generator $g_{\Phi}$
3:Initialize synthetic features $\mathbf{X}_{\mathcal{S}}$ and $g_{\Phi}$ parameters $\Phi$
4:for $k=1$ to $K$ do
5:  Initialize GNN parameters $\theta_{0}\sim P_{\theta_{0}}$
6:  for $t=0$ to T1T-1 do
7:   Sample a batch $\mathcal{T}^{\prime}$ from $\mathcal{T}$
8:   Compute real data gradient $\mathbf{g}_{\mathcal{T}}=\nabla_{\theta}\mathcal{L}_{\mathcal{T}^{\prime}}(\theta_{t})$
9:   Generate synthetic adjacency matrix $\mathbf{A}_{\mathcal{S}}=g_{\Phi}(\mathbf{X}_{\mathcal{S}})$ using Eq. (8)
10:   Compute synthetic loss $\mathcal{L}_{\mathcal{S}}(\theta_{t})$ on $\mathcal{S}=(\mathbf{X}_{\mathcal{S}},\mathbf{A}_{\mathcal{S}})$
11:   Compute synthetic model gradient $\mathbf{g}_{\mathcal{S}}=\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\theta_{t})$
12:   Compute synthetic feature gradient matrix $\mathbf{G}_{\mathcal{S}}=\nabla_{\mathbf{X}_{\mathcal{S}}}\mathcal{L}_{\mathcal{S}}(\theta_{t})$
13:// Component I: Gradient Decoupling
14:   Construct decoupled gradient matrix $\mathbf{G}^{\prime}_{\mathcal{S}}$ from $\mathbf{G}_{\mathcal{S}}$ using Eq. (10) (11)
15:// Component II: Structural Damping
16:   Compute structure regularization loss $\mathcal{R}_{\text{struct}}$ using Eq. (12)
17:// Meta-Update of Synthetic Data (Outer Loop)
18:   Compute gradient matching loss $\mathcal{L}_{\text{gm}}=D(\mathbf{g}_{\mathcal{S}},\mathbf{g}_{\mathcal{T}})$
19:   Compute total update loss $\mathcal{L}_{\text{update}}=\mathcal{L}_{\text{gm}}+\lambda\cdot\mathcal{R}_{\text{struct}}$
20:   Update $\mathbf{X}_{\mathcal{S}}$, $\Phi$ using $\nabla_{\mathbf{X}_{\mathcal{S}},\Phi}\mathcal{L}_{\text{update}}$ and learning rates $\eta_{\mathcal{S}}$, $\eta_{\Phi}$
21:// Update Model Parameters (Inner Loop Curriculum Step)
22:   Update model parameters $\theta_{t+1}\leftarrow\theta_{t}-\eta_{\theta}\mathbf{g}_{\mathcal{S}}$
23:  end for
24:end for
25:return Optimized synthetic features $\mathbf{X}_{\mathcal{S}}$ and generator parameters $\Phi$
4. Experiment
To empirically evaluate the effectiveness and generalization capability of our proposed SR-GM framework, we conducted a comprehensive set of experiments. This section is structured as follows: We begin by detailing the experimental setup, including datasets, evaluation protocols, and competing methods. Subsequently, we compare SR-GM against a series of representative baselines to demonstrate its superior performance. To further validate the robustness and architecture-agnostic nature of our framework, we then investigate its effectiveness when integrated with various GNN architectures. Furthermore, we perform experiments on SR-GM’s training convergence and modality correlation, aiming to explore its training efficiency and the impact of inter-modal conflict on multimodal graph condensation. Finally, we conduct extensive experiments on key hyperparameters—specifically, the damping coefficient $\lambda$ and the condensation ratio rr—thereby providing a thorough validation of SR-GM’s effectiveness on multimodal graph data. All the experiments are conducted on an NVIDIA RTX 3090 GPU.
Table 1. Statistics of datasets.
Name Nodes Edges Train/Val/Test
Ele-fashion 97,766 199,602 58,659/9,777/29,330
Goodreads-NC 685,294 7,235,084 406,689/67,805/203,428
Amazon-Sports 31,073 351,294 21,751/4,660/4,662
Amazon-Cloth 84,357 1,107,964 59,049/12,653/12,655
Table 2. Inductive performance on Ele-fashion , Goodreads-NC and Transductive performance on Amazon-Sports, Amazon-Cloth. Performance is reported as test accuracy (%) for inductive and AUC score for transductive. ”Original” denotes models trained on original graph. The best accuracies are bolded and the second best are underlined. $\triangle$(%) denotes the improvements of SR-GM upon the sub-optimal result.
Baselines Proposed
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCond SR-GM $\triangle$(%) Original
[…]
4.1. Experimental Setup
Datasets
We evaluate the condensation performance of the proposed framework on 4 multimodal datasets, i.e., Ele-fashion (ni2019justifying) and Goodreads-NC (wan2018item) for inductive setting and Amazon-Sports, Amazon-Cloth for transductive setting. Following (cai2024multimodal), we construct binary classification datasets for Amazon-Sports and Amazon-Cloth from the Amazon e-commerce platform [^1] https://www.amazon.com/; reviews with ratings $\geq$ 4 are labeled as positive, and those with ratings ¡ 2 as negative. Text and visual features for the Amazon-Sports and Amazon-Cloth datasets are encoded and aligned with ImageBind (girdhar2023imagebind), while CLIP (radford2021learning) is used for the Goodreads-NC and Ele-fashion datasets. The resulting features have dimensions of 1024 for Ele-fashion and Goodreads-NC, and 2048 for Amazon-Sports and Amazon-Cloth, and in all cases, the text and image modalities each constitute half of the total dimensionality. For the inductive setting, the test graph is not available during training. Dataset statistics are shown in Table 1.
Baselines
To demonstrate the effectiveness of SR-GM, we benchmark it against a comprehensive suite of baselines, which can be categorized into coreset selection and graph condensation methods. For coreset selection, we induce a subgraph from a representative node subset. We evaluate three selection strategies: Random, Herding (welling2009herding), and K-Center (sener2017active). For graph condensation, we employ three methods. GCond (jin2021graph) and MSGC (gao2023multiple) aim to preserve performance by matching GNN training gradients. In contrast, SGDD (yang2023does) focus on maintaining global graph properties by preserving the graph’s spectral characteristics and connectivity patterns.
Evaluation
We evaluate the quality of the condensed graphs by their ability to preserve performance on the downstream task of inductive node classification. For each baseline and our proposed SR-GM, we first generate a condensed graph containing $rN$ nodes from the original training graph of $N$ nodes, where $r\in(0,1)$ is the condensation ratio. The evaluation follows a two-stage protocol. In the training stage, a standard GNN model is trained exclusively on the small, condensed graph. Crucially, the original full graph is entirely discarded and inaccessible during this stage. In the subsequent test stage, the GNN model trained on the condensed graph is applied to the original, unseen test graph to perform inference on the test nodes. The primary metric for evaluation is the resulting test accuracy. To ensure robust conclusions, we repeat all experiments 5 times and report the mean accuracy and its variance.
4.2. Comparison with Baselines
In this subsection, we test the performance of a 2-layer GCN on the condensed graphs, and compare the proposed SR-GM with baselines. Table 2 reports node classification performance. As shown in Table 2, our proposed SR-GM achieves performance that is comparable to, and often superior to, that of the baseline methods across all datasets and condensation ratios. We make the following observations:
Superiority of SR-GM.:
Our proposed SR-GM significantly outperforms most baseline methods across all datasets and condensation ratios. For instance, on the Goodreads-NC dataset at a 0.025% ratio, SR-GM achieves 66.43% accuracy, creating a substantial performance margin of nearly 9 percentage points over the strongest baseline, GCond.
Stability and Robustness.:
Beyond superior accuracy, SR-GM demonstrates exceptional stability, especially in settings where other methods falter. As seen in Table 2, GCond exhibit unstable performance on datasets like Ele-fashion. SGDD and MSGC even fail on Goodreads-NC. In stark contrast, SR-GM consistently produces high-performing condensed graphs with low variance across all runs. We attribute this robustness to our optimization objective, which is designed to mitigate the gradient conflict between different modalities.
Impact of condensation Ratio.:
As expected, the accuracy of most methods improves as the condensation ratio ($r$) increases, since a larger condensed graph can store more information. However, it is crucial to note that the performance gap between SR-GM and the baselines is most pronounced at lower condensation ratios. For example, on Ele-fashion, the accuracy improvement of SR-GM over the next-best baseline is larger at $r=0.1\%$ than at $r=0.5\%$.
Figure 2. Comparison of condensation efficiency between our proposed SR-GM and the baseline GCond on the Ele-fashion and Goodreads-NC datasets. We can observe that SR-GM achieves comparable or superior peak performance to GCond with a faster convergence speed.
4.3. Condensation Efficiency
In this section, we empirically evaluate the condensation efficiency of SR-GM against GCond. This is achieved by analyzing the relationship between the performance of a node classification model and the number of training epochs spent optimizing the condensed dataset. Figure 2 illustrates this relationship, with each curve representing the performance trajectory and terminating upon reaching its peak.
As shown in Figure 2, SR-GM converges faster by 220 and 50 epochs on the Ele-fashion and Goodreads-NC datasets, respectively. Results demonstrates that SR-GM performances on par with, or even superior to GCond while maintaining a comparable convergence speed.
Table 3. Evaluation of cross-architecture generalizability. Test accuracy (%) of various GNN architectures trained on a graph condensed by SR-GM (with a GCN backbone).SAGE stands for GraphSAGE. Avg. stands for the average test accuracy.
Dataset Methods GCN MLP MMGCN GAT SAGE Avg.
Ele-fashion GCond 78.87 74.94 71.53 34.65 64.46 64.89
$r=0.1\%$ SR-GM 79.91 81.71 75.88 39.31 81.50 71.66
Goodreads-NC GCond 62.07 49.06 43.81 33.41 53.56 48.38
$r=0.05\%$ SR-GM 67.76 57.34 51.36 34.33 64.16 54.99
Table 4. Test accuracy (%) of SR-GM on the Ele-fashion datasets when using different backbone models for both condensation and evaluation. SAGE is an abbreviation for GraphSAGE. Original means training on original graph.
Backbone Ratio ($r$) SR-GM GCond Original
[…]
4.4. Cross-Architecture Generalizability
To evaluate the cross-architecture generalizability of our method, we condense a graph using SR-GM with a GCN backbone and then use this single condensed graph to train several different GNN architectures. As reported in Table 3, we test a diverse set of models, including GraphSAGE (hamilton2017inductive), GAT (velickovic2017graph), a multimodal GNN (MMGCN) (wei2019mmgcn), and a simple MLP (taud2017multilayer). The performance across these varied architectures demonstrates that the graphs condensed by SR-GM are not overfitted to the condensation model and possess good generalization capabilities.
4.5. Architectural Agnosticism of SR-GM
A key advantage of our SR-GM framework is its architectural agnosticism, meaning it is not tightly coupled to a specific GNN architecture. To demonstrate this flexibility, we conducted an experiment where we varied the underlying graph neural network used within the SR-GM condensation process.
We selected a diverse set of models for this evaluation, including popular GNNs such as GCN and GraphSAGE. Crucially, we also included a standard MLP as a stress test. The MLP is entirely graph-agnostic and processes only node features, allowing us to evaluate SR-GM’s ability to condense structural information even when the internal model does not perform message passing.
As shown in Table 4, we report the performance on the Ele-fashion dataset, where for each run, the same model (e.g., GCN) was used for both condensation and downstream evaluation. The results compellingly show that SR-GM consistently yields highly effective condensed graphs across all tested architectures, including the MLP. This strongly validates our claim that SR-GM’s effectiveness is not contingent on a particular GNN backbone, highlighting its robustness and wide applicability.
Figure 3. Test accuracy (%) for SR-GM and GCond using different input features: text-only, image-only, and fused text-image features on Ele-fashion and Goodreads-NC.
4.6. Ablation Study on the Impact of Multimodal Features
To investigate the impact of multimodal features on the performance of graph condensation, we conduct a comprehensive ablation study. In this experiment, we apply the baseline GCond method to graphs constructed with three different types of input features: text-only, image-only, and fused text-image features. We then compare its performance against our proposed SR-GM, which is implemented within the GCond framework to specifically address multimodal challenges. The node classification test accuracy on the Ele-fashion and Goodreads-NC datasets for these configurations is presented in Figure 3.
As illustrated in the figure, the introduction of fused multimodal data adversely affects the standard GCond method. We observe that the inherent conflict between different modalities leads to an average performance degradation of approximately 6% for GCond when compared to its performance on single-modality inputs. This highlights a significant challenge for existing gradient-matching-based condensation methods when applied to multimodal graphs. In contrast, our proposed SR-GM successfully mitigates this inter-modality gradient conflict during the condensation process. Consequently, SR-GM significantly outperforms GCond on the multimodal graph condensation task, achieving substantial accuracy improvements of 14% on Ele-fashion and 9% on Goodreads-NC. This result underscores the robustness and effectiveness of SR-GM in handling complex, multimodal graph data.
Figure 4. Heatmap of test accuracy for SR-GM on the Ele-fashion dataset under different settings of condensation rate ($r$) and regularization weight ($\lambda$). Darker shades correspond to higher accuracy, indicating better performance.
4.7. Parameter Analysis of Regularization Weight $\lambda$
A critical hyperparameter in SR-GM is the regularization weight, denoted by $\lambda$, which significantly influences the model’s performance. To investigate its impact on the graph condensation process, we conduct experiments on the Ele-fashion dataset under various settings of $\lambda$ and condensation ratio ($r$). The resulting test accuracies are visualized in a heatmap, as illustrated in Figure 4, where darker shades indicate higher accuracy.
From the results, we draw two primary observations. First, SR-GM demonstrates robust performance across most configurations, achieving high accuracy on the Ele-fashion dataset. A notable exception occurs at $r$ of 0.05%, where performance degrades for $\lambda$ values of 500 and 1000. Second, the findings suggest that while SR-GM is largely insensitive to high values of $\lambda$, an important correlation exists between $r$ and $\lambda$. Specifically, as $r$ increases (i.e., the number of nodes in the condensed graph grows), a correspondingly larger $\lambda$ is required to maintain optimal performance and prevent a significant decline in accuracy.
5. Conclusion
In this paper, we identify two fundamental challenges leading to the failure of current multimodal graph condensation methods: gradient conflicts arising from synthetic node features and the pathological amplification of gradient noise by graph structures. To address these issues, we propose a novel framework, Structurally-Regularized Gradient Matching (SR-GM), for condensing multimodal graph data. Specifically, SR-GM tackles the problem through a bi-level optimization formulation that incorporates two synergistic components. First, a gradient disentanglement mechanism resolves inter-modal conflicts at their source via orthogonal projection. Second, a structural damping regularizer, based on the Dirichlet energy of the gradient field, transforms the graph topology from a noise amplifier into an optimization stabilizer. Experimental results demonstrate the superiority of the proposed SR-GM in both efficiency and effectiveness. Furthermore, SR-GM exhibits strong generalization capabilities across various GNN architectures on multimodal graph-structured data.
Appendix A APPENDIX
Table 5. Inductive performance on Ele-fashion, Goodreads-NC and Transductive performance on Amazon-Sports, Amazon-Cloth using T5VIT encoding. Performance is reported as test accuracy (%) for inductive and AUC score for transductive. ”Original” denotes models trained on original graph. The best accuracies are bolded and the second best are underlined. $\triangle$(%) denotes the improvements of SR-GM upon the sub-optimal result.
Baselines Proposed
Dataset Ratio ($r$) Random Herding K-Center MSGC SGDD GCond SR-GM $\triangle$(%) Original
[…]
Table 6. SR-GM Ablation Study Results. D: gradient decoupling, R: structure regularization. $\Delta_{1}$: (SR-GM) - (SR-GM w/o D), $\Delta_{2}$: (SR-GM w/o D) - (SR-GM w/o D+R).
Dataset Ratio ($r$) Method CLIP (Aligned) T5VIT (Unaligned) CLIP Increments T5VIT Increments
[…]
A.1. Analysis of Feature Alignment Effects on Multimodal Graph Condensation
We conduct multimodal graph condensation experiments with unaligned feature encoding to simulate real-world scenarios with feature inconsistency, where textual features are encoded using T5 (raffel2020exploring) (512-dimensional) and visual features using ViT (dosovitskiy2020image) (512-dimensional). This heterogeneous encoding creates semantic gaps in the feature space, providing a challenging environment to validate SR-GM’s robustness. As shown in Table 5, SR-GM maintains superior performance across all datasets and condensation ratios under unaligned conditions, achieving improvements of 7.81% over the sub-optimal method on Goodreads-NC at 0.05% condensation ratio and 3.11% on Amazon-Sports at 0.3% condensation ratio, confirming the effectiveness of gradient decoupling in mitigating multimodal gradient conflicts.
Comparative analysis between Table 2 (aligned features) and Table 5 (unaligned features) reveals that aligned conditions yield higher original dataset accuracy, with Ele-fashion achieving 85.68% versus 84.47%, Goodreads-NC reaching 79.45% versus 76.93%, and Amazon-Cloth attaining 71.59% versus 68.31%, reinforcing the crucial role of feature alignment in multimodal graph learning. However, the radar charts in Figure 5 demonstrate SR-GM’s distinctive advantage: in both aligned and unaligned scenarios, it exhibits the smallest performance gap relative to original datasets, with Figure 5 particularly showing significantly smaller performance degradation compared to other condensation algorithms under unaligned conditions.
These findings highlight SR-GM’s unique capability to maintain excellent performance in unaligned environments through its innovative gradient decoupling and structural regularization mechanisms, substantially reducing dependency on perfect feature alignment. This strong adaptability to unaligned feature environments positions SR-GM as a powerful solution for practical multimodal graph condensation challenges, especially considering that achieving perfect feature alignment in real-world applications is often prohibitively expensive or practically infeasible.
Figure 5. Average performance comparison between feature aligned and feature misaligned
A.2. Ablation Study on Gradient Decoupling and Structure Regularization
The ablation study results fully validate the effectiveness of the SR-GM framework in addressing gradient conflict issues in multimodal graph condensation. In feature-aligned CLIP encoding scenarios, structure regularization (R) demonstrates significant advantages, particularly on the Goodreads-NC dataset where the $\Delta_{2}$ increments reach +7.56 to +8.88. This confirms our theoretical hypothesis regarding “structural damping” - by constraining the gradient field through Laplacian regularization, it effectively suppresses the propagation and amplification of local gradient conflicts within the graph structure. However, in high condensation ratio scenarios on the Ele-fashion dataset, the method relying solely on structure regularization ($\Delta_{1}=-1.18$) shows limitations, reflecting that in feature-aligned and data-sufficient environments, structural constraints alone may be insufficient to completely resolve the deep optimization problems caused by inter-modal semantic gaps.
In contrast, the core value of gradient decoupling (D) becomes prominent in feature-unaligned T5VIT encoding environments. The stable $\Delta_{1}$ increments of +2.58 to +4.00 on the Ele-fashion dataset directly confirm the crucial role of the gradient decoupling mechanism in handling conflicts among multimodal gradient signals. Particularly noteworthy is the negative $\Delta_{2}$ increment observed for structure regularization under T5VIT encoding, a phenomenon that profoundly reveals how traditional structural constraints might instead reinforce incorrect gradient propagation paths in feature-unaligned environments. Gradient decoupling, through orthogonal projection to eliminate inter-modal gradient contradictions, provides a cleaner optimization foundation for structure regularization.
These findings strongly support the dual-design philosophy of SR-GM: gradient decoupling resolves multimodal gradient conflicts at their source, while structure regularization suppresses conflict amplification along propagation paths. The synergistic effect of these two components enables SR-GM to maintain robust performance across diverse multimodal graph scenarios, providing reliable technical support for practical applications of multimodal graph condensation.